OpenStudy (cutiecomittee123):

PRE CALC HELP A hot bowl of soup cools according to Newtons Law of Cooling. Its temperature (in degrees Fahrenheit) at time t given by < T(t)=68+144e^0.04t. where t is given in minutes. What is the temperature of the soup after 15 minutes? thus far I wrote T(15)=68+144e^-0.04(15)

OpenStudy (toxicsugar22):

I dont know srry

OpenStudy (toxicsugar22):

i am not in calculus

OpenStudy (cutiecomittee123):

Darn it, thanks anyways

OpenStudy (16shuston):

im in precalc.. give just a sec

OpenStudy (cutiecomittee123):

Thanks:)

OpenStudy (16shuston):

ur doing everything correct so far.. ur next step is to multiply and add everything together

OpenStudy (16shuston):

basically plug all that u have into ur calculate.. and thats ur answer

OpenStudy (anonymous):

c. 125=68+144e^(-0.4t) (125-68)/144=0.395833333=e^(-.04t) -0.926762032=.04t t=23.169 After 23 minutes the soup is 125°F

OpenStudy (16shuston):

u dont have to do that, thats finding t, they r trying to find after 15 mins

OpenStudy (anonymous):

CAN SOMEONE HELP ON MY QUESTION NOW? I JUST ANSWERED UR QUESTION <33333333333

OpenStudy (anonymous):

HELP

OpenStudy (cutiecomittee123):

Okay thanks :)))

OpenStudy (cutiecomittee123):

@xoeindi give me a link to your question

OpenStudy (anonymous):

ok hold on

OpenStudy (cutiecomittee123):

Wait I just plugged in what I had to my calculator and I get 330.3 degrees farenheit, but the initial temperature is 212...Something is wrong here.

OpenStudy (cutiecomittee123):

@16shuston can you help??

OpenStudy (16shuston):

was ur calculator in degrees or radians?

OpenStudy (johnweldon1993):

It is just in the typo u there...it is -0.04 not the 0.04 as written \[\large T(t) = 68 + 144e^{-0.04(15)} = ?\]

jimthompson5910 (jim_thompson5910):

what kind of calculator do you have? how are you typing in the expression?

OpenStudy (johnweldon1993):

You will indeed get that 330 temperature if you do the 0.04 @cutiecomittee123 So make sure you are using the -0.04

OpenStudy (cutiecomittee123):

oh thanks!! gotcha